# Revisiting Monocular 3D Object Detection from Scene-Level Depth Retargeting to Instance-Level Spatial Refinement

### v1

@@ Abstract
Monocular 3D object detection is challenging due to the lack of accurate depth. However, existing depth-assisted solutions still exhibit inferior performance, whose reason is universally acknowledged as the unsatisfactory accuracy of monocular depth estimation models. In this paper, we revisit monocular 3D object detection from the depth perspective and formulate an additional issue as the limited 3D structure-aware capability of existing depth representations (e.g., depth one-hot encoding or depth distribution). To address this issue, we propose a novel depth-adapted monocular 3D object detection network, termed RD3D, that mainly comprises a Scene-Level Depth Retargeting (SDR) module and an Instance-Level Spatial Refinement (ISR) module. The former incorporates the scene-level perception of 3D structures, retargeting traditional depth representations to a new formulation: Depth Thickness Field. The latter refines the voxel spatial representation with the guidance of instances, eliminating the ambiguity of 3D occupation and thus improving detection accuracy. Extensive experiments on the KITTI and Waymo datasets demonstrate our superiority to existing state-of-the-art (SoTA) methods and the universality when equipped with different depth estimation models. The code will be available.
@@ 1 Introduction
3D object detection is critical in computer vision and has a wide range of applications, such as autonomous driving, video surveillance, etc. Early LiDAR-based methods [40, 14, 32] have achieved promising performance but suffer from high hardware costs. Although several exclusively multi-view methods [28, 8, 17, 15] have gained considerable success, they typically require multiple cameras and suffer from view inconsistency. Recently, there have been considerable advancements in monocular methods [27, 21, 19] and garnered substantial attention from both academic and industrial communities, which give credit to their low cost and ease of deployment.
@@
To address this issue, we propose RD3D, a novel depth-adapted monocular 3D detection network. Different from existing approaches [30, 27] that adapt Lift-Splat-Shoot (LSS) [28] framework and bird’s-eye-view (BEV) space, RD3D retargets traditional depth representations in a new formulation: Depth Thickness Field to adapt to 3D object detection. The depth thickness field can not only reduce the impact of erroneous depth estimation but also integrate the thickness property for each image pixel in the 3D space. It implicitly describes the occupation thickness along camera rays, producing sharper and more accurate features, as illustrated in Fig. 1. RD3D consists of two main components. First of all, a Scene-Level Depth Retargeting (SDR) module is presented to retarget traditional depth representations to the proposed depth thickness field. Compared with previous representations, it embeds clear 3D structure awareness, benefiting in distinguishing different instances and backgrounds. Then, we sample image features and depth thickness field to establish the voxel space with predefined grids. Second, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space by raising the oversight of individual instances. The network can focus more on the location and volume of instances rather than the whole 3D scene, eliminating the ambiguity of 3D occupation and thereby improving detection accuracy.
@@
  • We investigate an additional issue of depth-assisted methods as the limited 3D structure-aware capability of existing depth representations and present RD3D, a depth-adapted monocular 3D detection network to solve it.
  • To incorporate the scene-level perception of 3D structures, we propose a Scene-Level Depth Retargeting (SDR) module that retargets traditional depth representations to a new one: depth thickness field, yielding sharp spatial features beneficial for 3D detection.
  • To further improve accuracy, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space and eliminate the ambiguity of 3D occupation.
  • Compared with SoTA solutions, the proposed RD3D demonstrates superior accuracy and generalizability.
Figure 2: The overall framework of our proposed RD3D. The input image is first sent to the backbone to extract the features, and the depth map is transformed to depth one-hot encoding in parallel. The Scene-Level Depth Retargeting (SDR) module implicitly learns 3D scene structure features, retargeting depth one-hot encoding to depth thickness field (Sec. 3.3). The Instance-Level Spatial Refinement (ISR) module refines the voxel features with the guidance of instances, eliminating the ambiguity of 3D occupation (Sec. 3.4). Finally, the detector is applied to predict the 3D bounding boxes.
@@ 2.2 Monocular Depth Estimation
Monocular depth estimation aims to generate a single depth value for every pixel in an image. Neural network-based methods [36, 7, 31] often regress continuous depth by aggregating color and geometric structure information from an image. However, these methods are susceptible to overfitting to a single scenario, a common issue in monocular depth estimation. Recent methods [1, 34] have effectively mitigated the problem by creating large-scale relative depth datasets to learn relative relations. DAMv2 [41] unleashes the power of large-scale unlabeled data through semantic segmentation-assisted supervision and data augmentation. Metric3Dv2 [42] introduces a canonical camera transformation to resolve metric depth ambiguity across different camera setups, greatly improving depth accuracy and generalization. In our work, we utilize the depth map generated by the depth estimation model as a depth prior for the 3D detection model, guiding the generation of the depth thickness field representation.
@@
Representing features in 3D space, such as BEV grids, removes scale differences in the targets and makes it easier for the network to learn feature scale consistency. Many works [15, 17, 8, 30, 22] employ BEV representations and achieve great success. LSS [28] generates point clouds by learning a pixel-level depth distribution and projecting it to obtain BEV features. BEVDepth [15] supervises the depth distribution using a projection of the depth map from the point cloud, thus improving the 3D detection performance. CaDDN [30] first used BEV representations for monocular 3D detection, and our approach follows its basic architecture design. Additionally, occupancy as an explicit representation of 3D space has recently experienced a surge in research. TPVFormer [10] employs sparse 3D occupancy labels from LiDAR as the supervision to obtain 3D features. OccupancyM3D [27] introduces occupancy to monocular 3D detection, exploring feature encoding and representation in frustum and voxel space. Unlike the above methods, we concentrate on combining occupancy representation and 3D bounding box (bbox) to explore the mutual support of occupancy and monocular 3D detection tasks.
@@ 3 Methodology
Figure 3: Comparisons between different depth representations. The two purple curves indicate the confidence of the pixel on each depth bin, obtained from three different depth representations. The left curve represents a thick object, and the right one represents a thin object. The depth thickness field has the unique ability to represent the object thickness along the camera ray.
3.1 Framework Overview
As illustrated in Figure 2, our framework takes a 3-channel RGB image and a single-channel depth map (generated by a pre-trained model) as input and outputs the 3D bounding box parameters of the object in the image. Initially, the RGB image $\mathbf{I}\in\mathbb{R}^{W_{I}\times H_{I}\times 3}$ is passed through the backbone (DLA34 [43]) to extract image features, while the depth map $\mathbf{D}\in\mathbb{R}^{W_{I}\times H_{I}}$ is converted into depth one-hot encoding representation. Next, the extracted image features $\mathbf{F}\in\mathbb{R}^{W_{F}\times H_{F}\times C}$ and depth one-hot encoding $\Psi\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ are fed into the SDR module for further processing. Specifically, we employ the depth one-hot encoding representation to transform the extracted features into frustum features format $\mathbf{G}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Then the depth thickness field $\mathbf{D}_{thickness}\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ is adaptively learned under the supervision of our proposed soft-extended one-hot encoding. This depth thickness field is then used to refine the original image features and produce new frustum features $\mathbf{\tilde{G}}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Following this, grid sampling is applied to voxelize the 3D features. The voxel features, denoted as $\mathbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$, are subsequently fed into the ISR module. This module refines the voxel space representation by learning an occupancy map, which is guided by the enhanced occupancy labels. Finally, the refined voxel features $\mathbf{\tilde{V}}\in\mathbb{R}^{X\times Y\times Z\times C}$ are compressed into a bird’s-eye view (BEV) space and forwarded to the 3D detection head to output the results.
@@ 3.2 Depth Thickness Field
By applying Eq. 1 to each pixel in the depth map, we can obtain the depth one-hot encoding of the entire map, denoted as $\Psi$. This representation is sensitive to depth variations, aiding in the distinction of different objects. However, it heavily depends on the accuracy of the depth map. Moreover, this method generates a narrow response on the surface of the 3D object, limiting its ability to fully capture the three-dimensional structure.
@@ 3.3 Scene-Level Depth Retargeting
(a) Occupancy map from sparse point clouds.
(b) Occupancy map from sparse point clouds and 3D bboxes.
Figure 5: An slice example of occupancy map. Purple boxes indicate positive occupancy samples, with darker shades representing those generated from point clouds and lighter shades from 3D bounding boxes. Grey boxes indicate negative occupancy samples. (b) is more complete than (a).
@@
As previously mentioned, the depth thickness field is adaptively generated by the network. To ensure that it accurately reflects object thickness, we impose the following constraints. Firstly, we define a mask to extend the response range within the standard depth one-hot encoding:
@@ 3.4 Instance-Level Spatial Refinement
Figure 6: An slice example of occupancy map. The 3D bbox do not precisely conform to the object’s shape, resulting in false positive occupancy samples as the dark purple points.
@@
GUPNet [23] CVPR21 22.26 15.02 13.12 30.29 21.19 18.20 22.76 16.46 13.72
MonoFlex [45] CVPR21 19.94 13.89 12.07 28.23 19.75 16.89 19.94 13.89 12.07
CaDDN [30] CVPR21 19.17 13.41 11.46 27.94 18.91 17.19 23.57 16.31 13.84
@@
RD3D(Ours) - 32.02 21.19 18.80 42.67 28.92 25.89 36.17 26.76 23.67
Improvement v.s. second-best +1.84 +0.17 +0.64 +7.29 +2.97 +3.3 +3.99 +4.74 +3.83
@@
Occupancy Labels. To generate more adapted occupancy labels for 3D object detection, we propose leveraging 3D bounding boxes to assist sparse point clouds in representing object thickness, as illustrated in Figure 5(b). Similar to OccupancyM3D [27], we denote the 3D occupancy labels derived from sparse point clouds as $O_{point}^{*}\in\mathbb{R}^{X\times Y\times Z}$, where $X$, $Y$, and $Z$ are determined by the predefined voxel size and detection range.
@@
The ISR module enhances the perception of instances by adding occupancy labels from the 3D bounding boxes and then using it to refine the voxel spatial representation which is beneficial to improve detection accuracy.
@@ 4.1 Implementation Details
We employ PyTorch [25] for implementation. The network is trained on 3 NVIDIA 4090 (24G) GPUs. We use Adam [12] optimizer with an initial learning rate of 0.001 and employ the one-cycle learning rate policy [33]. On KITTI dataset [6], we train the model for 80 epochs with a total batch size of 12 and use [2, 46.8] $\times$ [-30.08, 30.08] $\times$ [-3, 1] (meter) for the range. On Waymo dataset [5], we downsample the input size to 640 $\times$ 960 and train 12 epochs with a detection range [2, 59.6] $\times$ [-25.6, 25.6] $\times$ [-2, 2] (meter). The voxel size is [0.16, 0.16, 0.16] (meter).
@@ 4.2 Benchmarks and Metrics
KITTI 3D. KITTI 3D is the most widely used benchmark [6] for 3D object detection, consisting of 7,481 training samples and 7,518 testing samples. Only annotations of the training set are made public. The training set is commonly divided into KITTI train with 3712 samples and KITTI val with 3769 samples following [3]. There are three difficulty levels, easy, moderate, and hard, defined by bounding box height and occlusion level. Following common practice, we use $AP_{3D}|_{R_{40}}$ and $AP_{BEV}|_{R_{40}}$ under IoU threshold of 0.7 to evaluate the performance.
Waymo. The Waymo Open Dataset [5] is a large dataset that contains 798 sequences for training and 202 sequences for validation. Following previous works [30], we report the performance on Waymo val set for the vehicle category using samples only from the front camera. Results on Waymo val are measured by officially evaluation of the mean average precision (mAP) and the mean average precision weighted by heading (mAPH).
@@ 4.3 Main Results
KITTI 3D. Table 1 shows the main results of RD3D on KITTI test and val sets. Our method achieves the best results under three-level difficulty settings in both $AP_{BEV}$ and $AP_{3D}$. On the test set, RD3D exceeds all existing methods and surpasses the second-best under three-level difficulties by +1.84/+0.17/+0.64 in $AP_{3D}$, and by +7.29/+2.97/+3.33 in $AP_{BEV}$. In particular, RD3D does not require the depth ground truth (GT) which is hard to obtain for supervision when training the detection model. Guided by accurate depth priors, our method transforms traditional depth representations into a more effective depth thickness field for 3D detection, resulting in a significant improvement over existing approaches.
Waymo. We also evaluate our method on Waymo Open Dataset (Waymo) [5]. As shown in Table 2, our method obtains promising results, significantly outperforming existing methods in 0-30m, 30-50m, and Overall. For example, under LEVEL 1 setting, RD3D outperforms the second best by 7.38/7.23 mAP/mAPH (18.04/17.79 vs. 10.66/10.56) and 9.78/9.75 mAP/mAPH (42.06/41.24 vs. 32.28/31.49) with IoU 0.7 and 0.5 criterions respectively; under LEVEL 2 setting, RD3D outperforms the second best by 6.74/6.60 mAP/mAPH (16.77/16.54 vs. 10.03/9.94) and 8.47/7.96 mAP/mAPH (39.20/38.44 vs. 30.73/30.48) with IoU 0.7 and 0.5 criterions respectively. This success can be attributed to the effective depth thickness field representation. For objects within [50m, $\infty$], our method performs relatively worse. This is because our method is voxel-based, which has a detection range limitation: [2, 59.6] (meters).

### v2

# Revisiting Monocular 3D Object Detection with Depth Thickness Field

@@ Abstract
Monocular 3D object detection is challenging due to the lack of accurate depth. However, existing depth-assisted solutions still exhibit inferior performance, whose reason is universally acknowledged as the unsatisfactory accuracy of monocular depth estimation models. In this paper, we revisit monocular 3D object detection from the depth perspective and formulate an additional issue as the limited 3D structure-aware capability of existing depth representations (e.g., depth one-hot encoding or depth distribution). To address this issue, we introduce a novel Depth Thickness Field approach to embed clear 3D structures of the scenes. Specifically, we present MonoDTF, a scene-to-instance depth-adapted network for monocular 3D object detection. The framework mainly comprises a Scene-Level Depth Retargeting (SDR) module and an Instance-Level Spatial Refinement (ISR) module. The former retargets traditional depth representations to the proposed depth thickness field, incorporating the scene-level perception of 3D structures. The latter refines the voxel space with the guidance of instances, enhancing the 3D instance-aware capability of the depth thickness field and thus improving detection accuracy. Extensive experiments on the KITTI and Waymo datasets demonstrate our superiority to existing state-of-the-art (SoTA) methods and the universality when equipped with different depth estimation models. The code will be available.
@@ 1 Introduction
3D object detection is critical in computer vision and has a wide range of applications, such as autonomous driving, robotic perception, etc. Although the methods based on LiDAR [44, 16, 35] and multi-view images [31, 9, 19, 17] have achieved promising performance, the heavy dependence on the expensive equipment [4, 26] restricts the application of these methods. Recently, there have been considerable advancements in monocular methods [30, 23, 21, 29, 42] and garnered substantial attention from both academic and industrial communities, which give credit to their low cost and ease of deployment.
@@
To address this issue, we propose MonoDTF, a novel scene-to-instance depth-adapted monocular 3D detection network. Different from existing approaches [33, 30] that adapt Lift-Splat-Shoot (LSS) [31] framework and bird’s-eye-view (BEV) space, MonoDTF retargets traditional depth representations in a new formulation: Depth Thickness Field (DTF) to adapt to 3D object detection. The DTF can reduce the impact of erroneous depth estimation and integrate the thickness property for each image pixel in the 3D space. It implicitly describes the thickness along camera rays, producing sharper and more accurate features, as illustrated in Fig. 1. MonoDTF consists of two main components. First of all, a Scene-Level Depth Retargeting (SDR) module is presented to retarget traditional depth representations to the proposed depth thickness field. Compared with previous representations, the DTF embeds clear 3D structure awareness, benefiting in distinguishing different instances and backgrounds. Then, we sample image features and depth thickness field to establish the voxel space with predefined grids. Second, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space by raising the oversight of individual instances. The network can focus more on the location and volume of instances than the whole 3D scene, enhancing the 3D instance-aware capability of the proposed depth thickness field and thereby improving detection accuracy.
@@
  • We investigate the issue of depth-assisted methods as the limited 3D structure-aware capability of existing depth representations and present a network (termed MonoDTF) with a novel depth representation, Depth Thickness Field, to solve it.
  • We design a Scene-Level Depth Retargeting (SDR) module that retargets traditional depth representations to the proposed Depth Thickness Field, incorporating the scene-level perception of 3D structures.
  • To enhance the 3D instance-aware capability of the depth thickness field, an Instance-Level Spatial Refinement (ISR) module is proposed to refine the voxel space and further improve detection accuracy.
  • Compared with SoTA solutions, the proposed MonoDTF demonstrates superior accuracy and generalizability.
Figure 2: The overall framework of our proposed MonoDTF. The input image is first sent to the backbone to extract the features and the depth map is transformed to depth one-hot encoding in parallel. The Scene-Level Depth Retargeting (SDR) module implicitly learns 3D scene structure features, retargeting depth one-hot encoding to depth thickness field (Sec. 3.3). The Instance-Level Spatial Refinement (ISR) module refines the voxel features with the guidance of instances, eliminating the ambiguity of 3D occupation (Sec. 3.4). Finally, the detector is applied to predict the 3D bounding boxes.
@@ 2.2 Monocular Depth Estimation
Monocular depth estimation aims to generate a single depth value for every pixel in an image. Neural network-based methods [39, 8, 34] often regress continuous depth by aggregating color and geometric structure information from an image. However, these methods are susceptible to overfitting to a single scenario, a common issue in monocular depth estimation. Recent methods [1, 37] have effectively mitigated the problem by creating large-scale relative depth datasets to learn relative relations. DAMv2 [45] unleashes the power of large-scale unlabeled data through semantic segmentation-assisted supervision and data augmentation. Metric3Dv2 [46] introduces a canonical camera transformation to resolve metric depth ambiguity across different camera setups, significantly improving depth accuracy and generalization. In our work, we utilize the depth map generated by the depth estimation model as a depth prior for the 3D detection model, guiding the generation of the depth thickness field representation.
@@
Representing features in 3D space, such as BEV grids, removes scale differences in the targets and makes it easier for the network to learn feature scale consistency. Many works [17, 19, 9, 33, 24] employ BEV representations and achieve great success. LSS [31] generates point clouds by learning a pixel-level depth distribution and projecting it to obtain BEV features. BEVDepth [17] supervises the depth distribution using a projection of the depth map from the point cloud, thus improving the 3D detection performance. CaDDN [33] first used BEV representations for monocular 3D detection, and our approach follows its basic architecture design. Additionally, occupancy as an explicit representation of 3D space has recently experienced a surge in research. TPVFormer [11] employs sparse 3D occupancy labels from LiDAR as the supervision to obtain 3D features. Unlike the above methods, we concentrate on combining occupancy representation and 3D bounding box (bbox) to explore the mutual support of occupancy and monocular 3D detection tasks.
@@ 3 Methodology
Figure 3: Comparisons between different depth representations. The two curves indicate the confidence of the pixel on each depth bin, obtained from three different depth representations respectively. The purple curve represents a thick object, and the claret one represents a thin object. The depth thickness field has the unique ability to represent the object thickness along the camera ray.
3.1 Preliminary and Overview
Task Definition. Given a single RGB image, monocular 3D object detection aims to recognize objects of interest and predict their 3D attributes, including 3D location $(x,y,z)$, dimension $(h,w,l)$, and orientation $\theta$. Existing approaches mainly focus on outdoor dynamic objects, with particular emphasis on the Car category [26].
Framework Overview. As illustrated in Figure 2, our framework takes a 3-channel RGB image and a single-channel depth map (generated by a pre-trained model) as input and outputs the 3D bounding box parameters of the object in the image. Initially, the RGB image $\mathbf{I}\in\mathbb{R}^{W_{I}\times H_{I}\times 3}$ is passed through the backbone (DLA34 [47]) to extract image features, while the depth map $\mathbf{D}\in\mathbb{R}^{W_{I}\times H_{I}}$ is converted into depth one-hot encoding representation. Next, the extracted image features $\mathbf{F}\in\mathbb{R}^{W_{F}\times H_{F}\times C}$ and depth one-hot encoding $\Psi\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ are fed into the SDR module for further processing. Specifically, we employ the depth one-hot encoding representation to transform the extracted features into frustum features format $\mathbf{G}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Then the depth thickness field $\mathbf{D}_{thickness}\in\mathbb{R}^{W_{F}\times H_{F}\times D}$ is adaptively learned under the supervision of our proposed soft-extended one-hot encoding. This depth thickness field is then used to refine the original image features and produce new frustum features $\mathbf{\tilde{G}}\in\mathbb{R}^{W_{F}\times H_{F}\times D\times C}$. Following this, grid sampling is applied to voxelize the 3D features. The voxel features, denoted as $\mathbf{V}\in\mathbb{R}^{X\times Y\times Z\times C}$, are subsequently fed into the ISR module. This module refines the voxel space representation by learning an occupancy map, which is guided by the enhanced occupancy labels. Finally, the refined voxel features $\mathbf{\tilde{V}}\in\mathbb{R}^{X\times Y\times Z\times C}$ are compressed into a bird’s-eye view (BEV) space and forwarded to the 3D detection head to output the results.
@@ 3.2 Depth Thickness Field
By applying Eq. 1 to each pixel in the depth map, we can obtain the depth one-hot encoding of the entire map, denoted as $\Psi$. This representation is sensitive to depth variations, aiding in distinguishing different objects. However, it heavily depends on the accuracy of the depth map. Moreover, this method generates a narrow response on the surface of the 3D object, limiting its ability to fully capture the three-dimensional structure.
@@ 3.3 Scene-Level Depth Retargeting
(a) Occupancy labels obtained from OccpancyM3D [30].
(b) Occupancy labels obtained from MonoDTF (Ours).
Figure 5: An slice example of occupancy map at a certain height. (a) OccupancyM3D takes voxels containing point clouds as positive samples (dark purple) and those passed through by camera ray and without point clouds as negative samples (gray). (b) Our method enhances (a) by incorporating voxels contained within the 3D boxes as positive samples (light purple).
@@
As previously mentioned, the depth thickness field is adaptively generated by the network. To ensure the DTF accurately reflects object thickness, we impose the following constraints. Firstly, we define a mask to extend the response range within the standard depth one-hot encoding:
@@ 3.4 Instance-Level Spatial Refinement
Figure 6: An slice example of occupancy map. Refining the 3D bounding box can reduce the False Positive Occupancy, thereby alleviating the ambiguity in the occupancy map.
@@
MonoDTF(Ours) - 32.02 21.19 18.80 42.67 28.92 25.89 36.17 26.76 23.67
Improvement v.s. second-best +1.84 +0.17 +0.64 +5.80 +2.97 +3.3 +3.99 +3.27 +3.83
@@
Occupancy Labels. To generate more complete and precise occupancy labels for 3D object detection, we propose leveraging 3D bounding boxes to assist sparse point clouds in representing object thickness. Firstly, similar to OccupancyM3D [30], we denote the 3D occupancy labels derived from sparse point clouds as $O_{point}^{*}\in\mathbb{R}^{X\times Y\times Z}$, where $X$, $Y$, and $Z$ are determined by the predefined voxel size and detection range, as illustrated in Figure 5(a). Then we denote an additional 3D occupancy label derived from 3D bounding boxes as $O_{bbox}\in\mathbb{R}^{X\times Y\times Z}$ to complement $O_{point}^{*}$, as shown in Figure 5(b).
@@
The ISR module enhances the perception of instances by adding occupancy labels from the 3D bounding boxes and then using it to refine the voxel spatial representation, which is beneficial in improving detection accuracy.
@@ 4.1 Implementation Details
We employ PyTorch [28] for implementation. The network is trained on 3 NVIDIA 4090 (24G) GPUs. We use Adam [14] optimizer with an initial learning rate of 0.001 and employ the one-cycle learning rate policy [36]. On KITTI dataset [7], we train the model for 80 epochs with a total batch size of 12 and use [2, 46.8] $\times$ [-30.08, 30.08] $\times$ [-3, 1] (meter) for the range of BEV space. On Waymo dataset [6], we downsample the input size to 640 $\times$ 960 and train 12 epochs with a detection range [2, 59.6] $\times$ [-25.6, 25.6] $\times$ [-2, 2] (meter). The voxel size for the KITTI and Waymo datasets is [0.16, 0.16, 0.16] (meter).
@@ 4.2 Benchmarks and Metrics
KITTI. KITTI is the most widely used benchmark [7] for 3D object detection, consisting of 7,481 training samples and 7,518 testing samples. Only annotations of the training set are made public. The training set is commonly divided into KITTI train with 3712 samples and KITTI val with 3769 samples following [3]. There are three difficulty levels, easy, moderate, and hard, defined by bounding box height and occlusion level. Following common practice, we use $AP_{3D}|_{R_{40}}$ and $AP_{BEV}|_{R_{40}}$ under IoU threshold of 0.7 to evaluate the performance.
Waymo. The Waymo Open Dataset [6] is a large dataset that contains 798 sequences for training and 202 sequences for validation. Following previous works [33, 30], we report the performance on Waymo val set for the vehicle category using samples only from the front camera. Results on Waymo val are measured by official evaluation of the mean average precision (mAP) and the mean average precision weighted by heading (mAPH).
@@ 4.3 Main Results
KITTI. Table 1 shows the main results of MonoDTF on KITTI test and val sets. Our method achieves the best results under three-level difficulty settings in $AP_{BEV}$ and $AP_{3D}$. On the test set, MonoDTF exceeds all existing methods and surpasses the second-best under three-level difficulties by +1.84/+0.17/+0.64 in $AP_{3D}$, and by +7.29/+2.97/+3.33 in $AP_{BEV}$. In particular, MonoDTF does not require the depth ground truth (GT) that is hard to obtain for supervision when training the detection model. Guided by accurate depth priors, our method transforms traditional depth representations into a more effective depth thickness field for 3D detection, resulting in a significant improvement over existing approaches.
Waymo. We also evaluate our method on Waymo Open Dataset (Waymo) [6]. As shown in Table 2, our method obtains promising results, significantly outperforming existing methods in “0-30m”, “30-50m”, and “Overall”. For example, under LEVEL 1 setting, MonoDTF outperforms the second best by +7.38/+7.23 and +9.78/+9.75 (mAP/mAPH) with IoU 0.7 and 0.5 criterions respectively; under LEVEL 2 setting, MonoDTF outperforms the second best by +6.74/+6.60 and +8.47/+7.96 (mAP/mAPH) with IoU 0.7 and 0.5 criterions respectively. This success can be attributed to the effective depth thickness field representation. For objects within [50m, $\infty$], our method performs relatively worse. This is because our method is BEV-based, which has a detection range limitation: [2, 59.6] (meters).